OpenStudy (anonymous):

Evaluate the function at each specified value of the independent variable and simplify. f(x)=x/x a.) f(4) b.)f(-4) c.) f(x-2)

OpenStudy (blurbendy):

f(4) = 4/4 = 1 f(-4) = -4/-4 = 1 f(x-2) = (x-2) / (x-2) = 1

OpenStudy (anonymous):

thanks! and part c.) also says \[\left\{ BLANK x< BLANK \right\}\] \[\left\{ BLANK X> BLANK \right\}\]

OpenStudy (blurbendy):

BLANK ?

OpenStudy (anonymous):

like the blank is the place for an answer

OpenStudy (blurbendy):

I don't understand. They all equal 1, so none of them is greater than or less than the other.

OpenStudy (anonymous):

this is why I hate pre calculus, makes NO sense